OpenStudy (anonymous):

How do I calculate the limit of 2^-x^2 as x approaches 0?

OpenStudy (zarkon):

this \[\lim_{x\to0}2^{-x^2}\] ?

OpenStudy (anonymous):

Yes. Sorry Im new at this.

OpenStudy (zarkon):

\[2^{-0^2}=2^0=1\]